The ratio of the magnitudes of the angular momentum of a projectile, at the highest point on trajectory and that at the point where it strikes the ground, about the point of projection is?

Grade:11

2 Answers

Chetan Mandayam Nayakar
13 years ago

Dear Varsha,

 

m((u2sin2(theta)/2g)*ucos(theta))/m((2u2sin(theta)cos(theta)/g)*usin(theta)=(1/2)/2=1/4

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Chetan Mandayam Nayakar – IIT Madras

vikas askiitian expert
13 years ago

angular momentam = mr*v

at highest point V = vcos@(i) &  r = R/2(i) + H(j)             

R is range & H is max height of projectile..

angular momentam=m(r*v)=mvcos@H (-k) ...............1

when the particle strikes the ground then its velocity v = vcos@(i)+vsin@(-j)

 r = R(i) + 0(j)

angular momentam=m(r*v)

                           =mvRsin@(-k)............2

ratio = eq1/eq2=cos@H/sin@R=1/4